OpenStudy (anonymous):

Which of the following is a solution to 5 square root of x minus 7 equals 10

OpenStudy (anonymous):

\[5\sqrt{x-7}=10 \rightarrow \sqrt{x-7}=2 \rightarrow x-7=4 \rightarrow x=11\]

OpenStudy (anonymous):

wow thankyou I really didn't get it!

OpenStudy (anonymous):

got it now?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

enjoy it, then